OpenStudy (rockinhood):

Brenton invested an average of $250 per month since age 39 in various securities for his retirement savings. His investments averaged a 6% annual rate of return until he retired at age 66. Given the same monthly investment and rate of return, how much more would Brenton have in his retirement savings had he started investing at age 25? Assume monthly compounding.

OpenStudy (rockinhood):

http://prntscr.com/bb6kx0

OpenStudy (retireed):

Use your equation given to you in your book or notes and calculate the amounts for both starting ages and subtract the larger one for the smaller one

OpenStudy (dumbcow):

sorry i misread the question, you need formula for sum of geometric series \[\rightarrow 250 (1 + (1+r) + (1+r)^2 + ...(1+r)^n)\] \[S(n) =250 \frac{(1+r)^n - 1}{r}\] r = monthly rate = 6% /12 = .5% n = num of months for each case

OpenStudy (rockinhood):

So how do I solve an equation of that size ??

OpenStudy (dumbcow):

you dont have to, you are given all the variables.... just put it into calculator

OpenStudy (dumbcow):

For 1st case. 66-39 = 27 yrs = 27*12 = 324 months ..... plug in n = 324 2nd case. 66-25 = 41 yrs = 41*12 = 492 months ...... plug in n = 492

OpenStudy (rockinhood):

I don't think I'm plugging it in right ?

OpenStudy (mathmate):

@rockinhood what did you get plugging the numbers in, 1. for 27 years, 2. for 41 years and 3. the difference?

OpenStudy (rockinhood):

decimals. 1. 5.66 2. 2.17 3. 3.49

OpenStudy (rockinhood):

@mathmate

hero (hero):

@dumbcow gave you more than enough to finish solving on your own imo.

hero (hero):

On a side note, you should invest in a good calculator--preferably ti-nspire.

OpenStudy (rockinhood):

Oops, found my problem. I was doing .5 instead of .005

OpenStudy (rockinhood):

Thanks for the.. encouragement ?? I guess

OpenStudy (rockinhood):

Yeah ended up with 5316.64 - 2016.37 _______________ 3,330.27

OpenStudy (mathmate):

From formula above: \(FV(P,i,n) =P \frac{(1+r)^n - 1}{r}\) FV=future value r=interest per month (in fractions, 0.06/12 for 6% per annum) n=number of months, (number of years * 12) If you are using a calculator, remember that all numbers in the numerator must be enclosed in parentheses. So please try the calculations again. Hint: the FV for 27 and 41 years both are in the hundreds of thousand dollars.

OpenStudy (mathmate):

You're close, but probably did something wrong with the $250 monthly payment. Please check the calculations.